# Quantum Algorithms in a Superposition of Spacetimes

### v1

@@ 4.2 Comparison to Standard Quantum Computation
As a brief explanation, the above fact follows by conditioning the execution of unitaries at each step, going (from left to write) on the elements of $\sigma$, and the unitary $U_{i}$ executes if and only if the current element of $\sigma$ is $i$. Finally, we put the above pieces together in order to simulate an execution of an OI oracle, using a regular quantum computer. The simulation of the OI oracle is without the amplification part.
@@
4.3 Sequentially Invertible Functions and Statistical Difference
As part of this work we define numerous new notions, with no dependence on any non-classical computing or information. Specifically, we define sequentially invertible functions, and a new computational problem, called the Sequentially Invertible Statistical Difference ($\mathsf{SISD}$) problem. We start with the definitions of sequential invertibility.
@@ 4.3 Sequentially Invertible Functions and Statistical Difference
For $r,t,\ell\in\mathbb{N}$, an $\left(r,t,\ell\right)$-invertible circuit sequence, is a sequence of $\ell$ pairs of randomized circuits $\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ such that for every $i\in[\ell],Y\in\{\rightarrow,\leftarrow\}$, the circuit $C_{i,Y}$ is of size $\leq t$, and such that,
  • There exists $\bar{k}\in\mathbb{N}$ such that all circuits have input and output size $\bar{k}$, and randomness bounded by $r$. Additionally, each pair of opposite-direction circuits use the same amount of randomness. Formally, for every $i\in[\ell]$ there exists $r_{i}\leq r$ such that
    $\forall Y\in\{\rightarrow,\leftarrow\}:C_{i,Y}:\{0,1\}^{\bar{k}}\times\{0,1\}^{r_{i}}\rightarrow\{0,1\}^{\bar{k}}\enspace.$
@@
    $\forall i\in[\ell],z\in\{0,1\}^{r_{i}},x\in\{0,1\}^{\bar{k}}:x=C_{i,\leftarrow}\left(C_{i,\rightarrow}\left(x;z\right);z\right)\enspace.$
@@
Definition 4.7 (Invertible circuit sequence sampling from $f$).
Let $r,t,\ell\in\mathbb{N}$, $\varepsilon\in[0,1]$, let $S:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ an $\left(r,t,\ell\right)$-invertible circuit sequence with input-output size $\bar{k}$, and let $f:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$ a function.
We say that $S$ samples $\varepsilon$-close to $f$ if, $\bar{k}=k$, and for the following distributions $D_{0}$, $D_{1}$ we have $\left\lVert D_{0}-D_{1}\right\rVert_{TV}\leq\varepsilon$,
$D_{0}=\Bigl{\{}f(z)\;|\;z\leftarrow\{0,1\}^{k^{\prime}}\Bigr{\}}$
$D_{1}=\biggl{\{}C_{\ell,\rightarrow}\left(\cdots C_{2,\rightarrow}\left(C_{1,\rightarrow}\left(0^{\bar{k}};z_{1}\right);z_{2}\right)\cdots;z_{\ell}\right)\;\Big{|}\;z_{1}\leftarrow\{0,1\}^{r_{1}},\cdots,z_{\ell}\leftarrow\{0,1\}^{r_{\ell}}\biggr{\}}\enspace.$
If $\varepsilon=0$, we simply say that $S$ samples from $f$.
Definition 4.8 ($\left(r,t,\ell,\varepsilon\right)$-sequentially invertible function).
Let $f:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$ a function, let $r,t,\ell\in\mathbb{N}$ and let $\varepsilon\in[0,1]$. We say that $f$ is $\left(r,t,\ell,\varepsilon\right)$-sequentially invertible if there exists an $\left(r,t,\ell\right)$-invertible circuit sequence $S$ that samples $\varepsilon$-close to $f$.
Definition 4.9 ($\left(r(n),t(n),\ell(n),\varepsilon(n)\right)$-sequentially invertible function family).
Let $f=\{f_{n}\}_{n\in\mathbb{N}}$ a family of functions: For every $n\in\mathbb{N}$, $f_{n}:\{0,1\}^{k^{\prime}(n)}\rightarrow\{0,1\}^{k(n)}$. Let $r,t,\ell:\mathbb{N}\rightarrow\mathbb{N}$ and let $\varepsilon:\mathbb{N}\rightarrow[0,1]$. We say that the family of functions $f$ is $\left(r(n),t(n),\ell(n),\varepsilon(n)\right)$-sequentially invertible if for every $n\in\mathbb{N}$, there exists an $\left(r(n),t(n),\ell(n)\right)$-invertible circuit sequence $S_{n}$ that samples $\varepsilon(n)$-close to $f_{n}$.
Note that when considering sequential invertibility, we can interchange between considering the sequential invertibility of functions and circuits.
Remark 4.2 (Interchanging between sequentially invertible circuits and functions).
While we define sequential invertibility for functions $f:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$, we sometimes refer to sequential invertibility for circuits $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$. This extension is natural, as circuits are in particular functions – a circuit contains not only the information of the function it computes, but a way to compute the function, using the standard set of classical gates $\{NOT,OR,AND\}$.
@@
where for each $b\in\{0,1\}$, $\left(C^{b}_{i,\rightarrow},C^{b}_{i,\leftarrow}\right)_{i\in[\ell]}$ is an $r(n)$-invertible circuit sequence, where $n$ is the input size,
$n:=\bigg{|}\left(\left(C^{0}_{i,\rightarrow},C^{0}_{i,\leftarrow}\right)_{i\in[\ell]},\left(C^{1}_{i,\rightarrow},C^{1}_{i,\leftarrow}\right)_{i\in[\ell]}\right)\bigg{|}\enspace,$
and such that both sequences have the same input and output size $k$, that is, for every $i\in[\ell]$, $b\in\{0,1\}$, $Y\in\{\rightarrow,\leftarrow\}$,
$C^{b}_{i,Y}:\{0,1\}^{k}\times\{0,1\}^{r^{b}_{i}(n)}\rightarrow\{0,1\}^{k}\enspace.$
@@
[…]
Note that in the above definition, while we consider sequentially invertible circuit sequences as the input to the problem, we focus only on $r$-invertibility, and ignore the parameters $t\in\mathbb{N}$, $\ell\in\mathbb{N}$, $\varepsilon\in[0,1]$ from the original definition. The reason is that all parameters, except $r$, are implicit in the problem input and problem definition. More precisely, the parameters $t,\ell$ are meaningless as the problem input is the circuit sequence (which defines $t,\ell$), and the total variation distance parameter $\varepsilon$ is ignored, as there are different total variation distance demands from the circuit sequences.
@@ 5 $\mathsf{SISD}_{\mathsf{poly},O\left(\log(n)\right)}\in\textbf{BQP}^{\textbf{OI}}$
In this section we prove that a quantum computer with access to an order interference (OI) oracle is able to solve $\mathsf{SISD}_{\mathsf{poly},O\left(\log(n)\right)}$ in polynomial time. For the result in this section, a uniform computable OI oracle (Definition 4.2) suffices.
@@
Next, we claim that (a factorial scaling of) a choice interference state can be expressed with an order interference state.
Lemma 5.1 (Norm of simulated choice interference).
Let $\{U_{i}\}_{i\in[m]}$ a set of $m\in\mathbb{N}$ unitary circuits, all operating on the same number of qubits $n\in\mathbb{N}$. Let $|{\psi}\rangle$ an $n$-qubit quantum state.
@@
Then,
$\textbf{OI}\left(\{\tilde{U}_{i}\}_{i\in[m]},|{\psi}\rangle|{0^{\lceil\log\left(m\right)\rceil}}\rangle\right)=\left(m-1\right)!\cdot\left(\textbf{OI}_{c}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\otimes|{m\text{ mod }2^{\lceil\log(m)\rceil}}\rangle\right)\enspace.$
@@
Together, this means that the OI state is,
@@
$:=\left(m-1\right)!\cdot\left(\textbf{OI}_{c}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\otimes|{m\text{ mod }2^{\lceil\log(m)\rceil}}\rangle\right)\enspace.$
Finally, the above implies the below corollary lemma, which says that an OI oracle can be used in order to simulate oracle access to choice interference with a factorial reduction in needed complexity, for successfully executing the transformation.
Lemma 5.2 (Choice interference oracle from order interference oracle).
Let $\mathcal{O}_{\textbf{OI}}\left(\cdot,\cdot,\cdot\right)$ a uniform order interference oracle as in Definition 4.2. There is a quantum procedure $Q_{CI}$ with oracle access to $\mathcal{O}_{\textbf{OI}}$ and the following input: $\{U_{i}\}_{i\in[m]}$ a set of $m\in\mathbb{N}$ unitary circuits, all operating on the same number of qubits $n\in\mathbb{N}$, $|{\psi}\rangle$ an $n$-qubit quantum state, and $\lambda\in\mathbb{N}$.
$Q_{CI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle,\lambda\right)$ executes in time $O\left(\lambda+\log\left(m\right)\cdot\left(\sum_{i\in[m]}|U_{i}|\right)\right)$ and with probability at least
$\frac{\frac{\left\lVert\textbf{OI}_{c}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m}}{\frac{\left\lVert\textbf{OI}_{c}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m}+\frac{1}{\lambda}}\enspace,$
outputs the normalization of the choice interference state $\textbf{OI}_{c}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)$.
@@
There exists $Q^{\textbf{OI}}\left(\cdot,\cdot\right)$ a quantum algorithm with oracle access to a uniform OI oracle (as in Definition 4.2), that gets as input (1) an amplification parameter $\delta\in\mathbb{N}$ and (2) an $r$-invertible circuit sequence $C:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ with output size $k$ of the circuits, and the total description size of the pair of sequences is denoted with $n$.
The algorithm $Q^{\textbf{OI}}$ executes in time $O\left(\delta\cdot\mathsf{poly}(n)\cdot 2^{r}\right)$ for some polynomial $\mathsf{poly}(\cdot)$, and outputs a success/fail bit $b\in\{0,1\}$ together with a $k$-qubit quantum state. The algorithm succeeds (outputs $b=1$) with probability $\geq 1-\frac{1}{\delta}$ and along with the bit outputs the (normalization of the) output distribution state of the circuit sequence:
@@
Next, observe that for every $i\in[\ell]$, for every $k$-qubit state $|{\phi}\rangle$ such that all of its amplitudes are real and non-negative, the vector $\sum_{z_{i}\in\{0,1\}^{r_{i}}}U_{i,z_{i}}\cdot|{\phi}\rangle$ has only real and non-negative amplitudes. This follows because for every $i\in[\ell]$, the unitaries in $\{U_{iz_{i}}\}_{z_{i}\in\{0,1\}^{r_{i}}}$ all map classical to classical states. It follows that the normalization of the choice interference state $\sum_{z_{i}\in\{0,1\}^{r_{i}}}U_{i,z_{i}}\cdot|{\phi}\rangle$ is a quantum state with only real and non-negative amplitudes.
@@
It thus follows that for every $k$-qubit state $|{\phi}\rangle$ such that all of its amplitudes are real and positive, we have,
@@
[…]

### v2

@@ 4.1 Quantum Computation with Computable Order Interference
Remark 4.2 (Stronger versions of the OI oracle).
The definition of the success probability we use in Definitions 4.2, 4.4 is restrictive in two regards. First, the non-amplifiable component,
$\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{\sigma\in S_{m}}\alpha_{x,\sigma}|}{\sum_{x\in\{0,1\}^{n}}\sum_{\sigma\in S_{m}}|\alpha_{x,\sigma}|}$
is intended to restrict the success probability, proportionally to how much the phases of states in different spacetimes (or, unitary evolution orders) are misaligned, as we do not assume a particular destructive interference dynamics between the same states from different spacetimes. The below Definition 4.6 is equivalent to the original definition of the OI oracle, without the phase alignment restriction. This definition was used as the main definition for the OI oracle in an earlier version of this work.
Second, the amplifiable norm component
is intended to restrict the ability to arbitrarily execute OI transformation, ignoring the norm of the OI state to-be-generated. The below Definition 4.7 is equivalent to the original definition of the OI oracle, without the norm amplification restriction. Note that in the case of Definition 4.7, the OI oracle’s input consists of one less parameter, as the amplification parameter $\lambda$ is dropped.
As noted in [Aar, Car] following an earlier version of this work, a quantum computer with access to a strengthened OI oracle as in Definition 4.6 allows for a polynomial time solution for the entire complexity class NP (and even QCMA). We leave it as an open problem what is the computational power of quantum computers with access to oracles as in Definitions 4.6, 4.7.
Definition 4.6 (Computable Order Interference Oracle without Phase Alignment).
A computable order interference oracle with no phase alignment, denoted $\mathcal{O}_{\textbf{OI}}^{\omega}$, is an oracle with input quadruple ($\left(\{U_{i}\}_{i\in[m]},R,U_{O},\lambda\right)$, such that,
  • $R$ is an $n$-qubit quantum register,
  • and $\lambda\in\mathbb{N}$ is a natural number.
$\sum_{y\in\{0,1\}^{n^{\prime}}}\alpha_{y}\cdot|{\psi_{y}}\rangle_{R}|{y}\rangle_{R^{\prime}}\enspace.$
$\min_{y\in\{0,1\}^{n^{\prime}}}\left(\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle,U_{O}\right)\right\rVert}{\sqrt{m!}}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle,U_{O}\right)\right\rVert}{\sqrt{m!}}+\frac{1}{\lambda}}\right)\enspace.$
$\sum_{y\in\{0,1\}^{n^{\prime}}}\alpha_{y}\cdot\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle,U_{O}\right)_{R}|{y}\rangle_{R^{\prime}}\enspace,$
- The unitaries $\{U_{i}\}_{i\in[m]}$ can also be given in the form of controlled oracle access. In that case, the complexity $|U_{i}|$ for $i\in[m]$ is the complexity of making an oracle call to the $(n+1)$-qubit unitary which is the controlled $U_{i}$.
Definition 4.7 (Computable Order Interference Oracle without Norm Amplification).
A computable order interference oracle without norm amplification, denoted $\mathcal{O}_{\left\lVert\textbf{OI}\right\rVert}$, is an oracle with input triplet ($\left(\{U_{i}\}_{i\in[m]},R,U_{O}\right)$, such that,
  • $R$ is an $n$-qubit quantum register,
Given valid input ($\left(\{U_{i}\}_{i\in[m]},R,U_{O}\right)$, let $R^{\prime}$ a $n^{\prime}$-qubit register such that the joint system $\left(R,R^{\prime}\right)$ is in a pure state, written as,
$\sum_{y\in\{0,1\}^{n^{\prime}}}\alpha_{y}\cdot|{\psi_{y}}\rangle_{R}|{y}\rangle_{R^{\prime}}\enspace.$
The oracle call takes time complexity $|U_{O}|+\sum_{i\in[m]}|U_{i}|$. At the end of the oracle execution it returns a success/failure bit $b\in\{0,1\}$, where the success probability is
$\min_{y\in\{0,1\}^{n^{\prime}}}\left(\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{\sigma\in S_{m}}\alpha_{y,x,\sigma}|}{\sum_{x\in\{0,1\}^{n}}\sum_{\sigma\in S_{m}}|\alpha_{y,x,\sigma}|}\right)\enspace,$
where for $y\in\{0,1\}^{n^{\prime}}$, $\sigma\in S_{m}$ we write
$\prod_{i\in[m]}U_{\sigma^{-1}\left(i\right)}\cdot|{\psi_{y}}\rangle:=\sum_{x\in\{0,1\}^{n}}\alpha_{y,x,\sigma}|{x}\rangle\enspace.$
$\sum_{y\in\{0,1\}^{n^{\prime}}}\alpha_{y}\cdot\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi_{y}}\rangle,U_{O}\right)_{R}|{y}\rangle_{R^{\prime}}\enspace,$
- The unitaries $\{U_{i}\}_{i\in[m]}$ can also be given in the form of controlled oracle access. In that case, the complexity $|U_{i}|$ for $i\in[m]$ is the complexity of making an oracle call to the $(n+1)$-qubit unitary which is the controlled $U_{i}$.
@@ 4.2 Comparison to Standard Quantum Computation
As a brief explanation, the above fact follows by conditioning the execution of unitaries at each step, going (from left to right) on the elements of $\sigma$, and the unitary $U_{i}$ executes if and only if the current element of $\sigma$ is $i$. Finally, we put the above pieces together in order to simulate an execution of an OI oracle, using a regular quantum computer. The simulation of the OI oracle is without the amplification part.
@@
4.3 Sequentially Invertible Distributions and Statistical Difference
As part of this work we define numerous new notions, with no dependence on any non-classical computing or information. Specifically, we define sequentially invertible distributions, and a new computational problem, called the Sequentially Invertible Statistical Difference ($\mathsf{SISD}$) problem. We start with the definitions of sequential invertibility.
@@ 4.3 Sequentially Invertible Distributions and Statistical Difference
For $r,t,\ell,k\in\mathbb{N}$, an $\left(r,t,\ell\right)$-invertible circuit sequence on $k$ bits, is a sequence of $\ell$ pairs of randomized circuits $\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ such that for every $i\in[\ell]$,
  • For every $Y\in\{\rightarrow,\leftarrow\}$, the circuit $C_{i,Y}$ is of size at most $t$, uses $r_{i}\leq r$ random bits and maps from $k$ to $k$ bits, that is,
    $C_{i,Y}:\{0,1\}^{k}\times\{0,1\}^{r_{i}}\rightarrow\{0,1\}^{k}\enspace.$
@@
    $\forall z\in\{0,1\}^{r_{i}},x\in\{0,1\}^{k}:x=C_{i,\leftarrow}\left(C_{i,\rightarrow}\left(x;z\right);z\right)\enspace.$
    Note that this also implies that for each hard-wired randomness, the circuits act as permutations on the set $\{0,1\}^{k}$.
@@
Definition 4.9 (Invertible circuit sequence sampling from $D$).
Let $r,t,\ell,k\in\mathbb{N}$, $\varepsilon\in[0,1]$, let $S:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ an $\left(r,t,\ell\right)$-invertible circuit sequence on $k$ bits, and let $D$ a distribution on $\{0,1\}^{k}$.
We say that $S$ samples $\varepsilon$-close to $D$ if for the following distribution $D_{S}$ we have $\left\lVert D_{S}-D\right\rVert_{TV}\leq\varepsilon$,
$D_{S}=\biggl{\{}C_{\ell,\rightarrow}\left(\cdots C_{2,\rightarrow}\left(C_{1,\rightarrow}\left(0^{k};z_{1}\right);z_{2}\right)\cdots;z_{\ell}\right)\;\Big{|}\;z_{1}\leftarrow\{0,1\}^{r_{1}},\cdots,z_{\ell}\leftarrow\{0,1\}^{r_{\ell}}\biggr{\}}\enspace.$
If $\varepsilon=0$, we simply say that $S$ samples from $D$.
Definition 4.10 ($\left(r,t,\ell,\varepsilon\right)$-sequentially invertible distribution).
Let $D$ a distribution, let $r,t,\ell\in\mathbb{N}$ and let $\varepsilon\in[0,1]$. We say that $D$ is $\left(r,t,\ell,\varepsilon\right)$-sequentially invertible if there exists an $\left(r,t,\ell\right)$-invertible circuit sequence $S$ that samples $\varepsilon$-close to $D$.
Definition 4.11 ($\left(r(n),t(n),\ell(n),\varepsilon(n)\right)$-sequentially invertible distribution family).
Let $D=\{D_{n}\}_{n\in\mathbb{N}}$ a family of distributions. Let $r,t,\ell:\mathbb{N}\rightarrow\mathbb{N}$ functions and let $\varepsilon:\mathbb{N}\rightarrow[0,1]$. We say that the family of distributions $D$ is $\left(r(n),t(n),\ell(n),\varepsilon(n)\right)$-sequentially invertible if for every $n\in\mathbb{N}$, there exists an $\left(r(n),t(n),\ell(n)\right)$-invertible circuit sequence $S_{n}$ that samples $\varepsilon(n)$-close to $D_{n}$.
Note that when considering sequential invertibility, we can interchange between considering the sequential invertibility of distributions and circuits.
Remark 4.3 (Interchanging between sequentially invertible circuits and distributions).
While we define sequential invertibility for a distribution $D$ over a set $\{0,1\}^{k}$, we sometimes refer to sequential invertibility for circuits $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$. This extension is natural, as circuits are in particular distributions – we can consider the output distributions a circuit $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$ given a uniformly random $k^{\prime}$-bit string. Additionally, a circuit contains not only the information of the distribution it computes, but a way to sample from the distribution, using the standard set of classical gates $\{NOT,OR,AND\}$.
@@
where for each $b\in\{0,1\}$, $\left(C^{b}_{i,\rightarrow},C^{b}_{i,\leftarrow}\right)_{i\in[\ell]}$ is an $r(n)$-invertible circuit sequence on some (identical) number of bits $k$ (as in Definition 4.8), where $n$ is the input size,
$n:=\bigg{|}\left(\left(C^{0}_{i,\rightarrow},C^{0}_{i,\leftarrow}\right)_{i\in[\ell]},\left(C^{1}_{i,\rightarrow},C^{1}_{i,\leftarrow}\right)_{i\in[\ell]}\right)\bigg{|}\enspace.$
@@
  • $\prod_{\text{YES}}$ is the set of pairs of sequences such that $\left\lVert D_{0}-D_{1}\right\rVert_{TV}\leq a(n)$,
  • $\prod_{\text{NO}}$ is the set of pairs of sequences such that $\left\lVert D_{0}-D_{1}\right\rVert_{TV}>b(n)$.
Note that in the above definition, while we consider sequentially invertible sampler sequences as the input to the problem, we focus only on $r$-invertibility, and ignore the parameters $t\in\mathbb{N}$, $\ell\in\mathbb{N}$, $\varepsilon\in[0,1]$ from the original definition. The reason is that all parameters, except $r$, are implicit in the problem input and problem definition. More precisely, the parameters $t,\ell$ are meaningless as the problem input is the sampler sequence (which defines $t,\ell$), and the total variation distance bound parameter $\varepsilon$ is ignored, as the total variation distance demands depend on $a(n)$ and $b(n)$.
@@ 5 $\mathsf{SISD}_{\mathsf{poly},O\left(\log(n)\right)}\in\textbf{BQP}^{\textbf{OI}}$
In this section we prove that a quantum computer with access to an order interference (OI) oracle is able to solve $\mathsf{SISD}_{\mathsf{poly},O\left(\log(n)\right)}$ in polynomial time. For the result in this section, a uniform computable OI oracle (Definition 4.2) suffices, and more so, the simplified Corollary 4.1 of the definition.
@@
The next lemma says that an OI oracle can be used in order to simulate oracle access to choice interference with a factorial reduction in needed complexity, for successfully executing the transformation.
Lemma 5.1 (Choice interference from order interference).
Let $\mathcal{O}_{\textbf{OI}}\left(\cdot,\cdot,\cdot\right)$ a uniform order interference oracle as in Definition 4.2. We define the procedure $Q_{CI}$, which is a quantum algorithm with oracle access to $\mathcal{O}_{\textbf{OI}}$. The procedure’s input is triplet ($\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle,\lambda\right)$ as in the case of the oracle $\mathcal{O}_{\textbf{OI}}$.
@@
$Q_{CI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle,\lambda\right)$ executes the oracle call
$\mathcal{O}_{\textbf{OI}}\left(\{\tilde{U}_{i}\}_{i\in[m]},|{\psi}\rangle|{0^{\lceil\log\left(m\right)\rceil}}\rangle,\lambda\right)\enspace,$
returns the success/fail bit of the oracle’s output and the output state, after tracing out the rightmost $\lceil\log\left(m\right)\rceil$ qubits. The procedure $Q_{CI}$ executes in time $O\left(\lambda+\log\left(m\right)\cdot\left(\sum_{i\in[m]}|U_{i}|\right)\right)$ and with probability at least
$\left(\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{i\in[m]}\alpha_{x,i}|}{\sum_{x\in\{0,1\}^{n}}\sum_{i\in[m]}|\alpha_{x,i}|}\right)\frac{\frac{\left\lVert\textbf{OI}_{c}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m}}{\frac{\left\lVert\textbf{OI}_{c}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m}+\frac{1}{\lambda}}\enspace,$
outputs the normalization of the choice interference state $\textbf{OI}_{c}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)$, where for $i\in[m]$, we define $U_{i}\cdot|{\psi}\rangle:=\sum_{x\in\{0,1\}^{n}}\alpha_{x,i}\cdot|{x}\rangle$.
@@
For $\sigma\in S_{m}$ we write
$\prod_{i\in[m]}U_{\sigma^{-1}\left(i\right)}\cdot|{\psi}\rangle:=\sum_{x\in\{0,1\}^{n}}\alpha_{x,\sigma}|{x}\rangle\enspace,$
and for $i\in[m]$, we define $U_{i}\cdot|{\psi}\rangle:=\sum_{x\in\{0,1\}^{n}}\alpha_{x,i}\cdot|{x}\rangle$. The above observations imply
$\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{\sigma\in S_{m}}\alpha_{x,\sigma}|}{\sum_{x\in\{0,1\}^{n}}\sum_{\sigma\in S_{m}}|\alpha_{x,\sigma}|}=\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{i\in[m]}\alpha_{x,i}|}{\sum_{x\in\{0,1\}^{n}}\sum_{i\in[m]}|\alpha_{x,i}|}\enspace.$
Also, the above observations imply,
@@
$:=\left(m-1\right)!\cdot\left(\textbf{OI}_{c}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\otimes|{m\text{ mod }2^{\lceil\log(m)\rceil}}\rangle\right)\enspace,$
which conclude our proof. ∎
@@
There exists $Q^{\textbf{OI}}\left(\cdot,\cdot\right)$ a quantum algorithm with oracle access to a uniform OI oracle (as in Definition 4.2), that gets as input (1) an amplification parameter $\delta\in\mathbb{N}$ and (2) an $r$-invertible circuit sequence $C:=\left(C_{i,\rightarrow},C_{i,\leftarrow}\right)_{i\in[\ell]}$ on $k$ bit strings, such that the total description size of the sequence is denoted with $n$.
The algorithm $Q^{\textbf{OI}}$ executes in time $O\left(\delta\cdot\mathsf{poly}(n)\cdot 2^{r}\right)$ for some polynomial $\mathsf{poly}(\cdot)$, and outputs a success/fail bit $b\in\{0,1\}$ together with a $k$-qubit quantum state $|{\phi}\rangle$. The algorithm succeeds (outputs $b=1$) with probability $\geq 1-\frac{1}{\delta}$ and in that case the state $|{\phi}\rangle$ is the (normalization of the) output distribution state of the circuit sequence:
@@
Next, observe that for every $i\in[\ell]$, for every $k$-qubit state $|{\phi}\rangle$ such that all of its amplitudes are real and non-negative, the vector $\sum_{z_{i}\in\{0,1\}^{r_{i}}}U_{i,z_{i}}\cdot|{\phi}\rangle$ has only real and non-negative amplitudes. This follows because for every $i\in[\ell]$, the unitaries in $\{U_{i,z_{i}}\}_{z_{i}\in\{0,1\}^{r_{i}}}$ all map classical to classical states. It follows that the normalization of the choice interference state $\sum_{z_{i}\in\{0,1\}^{r_{i}}}U_{i,z_{i}}\cdot|{\phi}\rangle$ is a quantum state with only real and non-negative amplitudes.
@@
It thus follows that for every $k$-qubit state $|{\phi}\rangle$ such that all of its amplitudes are real and non-negative, we have,
@@
Let $\mathsf{SISD}_{a,b,r}:=\left(\prod_{\text{YES}},\prod_{\text{NO}}\right)$ an $\mathsf{SISD}$ promise problem, such that $\mathsf{SISD}_{a,b,r}\in\mathsf{SISD}_{\mathsf{poly},\,O\left(\log(n)\right)}$. Then, there is $M^{\textbf{OI}}$ a QPT algorithm with oracle access to a computable uniform order interference oracle $\mathcal{O}_{\textbf{OI}}$ (Definition 4.2), that for every $x\in\prod_{\text{YES}}$, $M^{\textbf{OI}}(x)=1$ with probability at least $p(x)$ and for every $x\in\prod_{\text{NO}}$, $M^{\textbf{OI}}(x)=0$ with probability at least $p(x)$, such that
$\forall x\in\left(\prod_{\text{YES}}\cup\prod_{\text{NO}}\right):p(x)\geq 1-2^{-\mathsf{poly}\left(|x|\right)}\enspace,$